OpenStudy (anonymous):

Simplify the expression below and write it as a single logarithm: 3log(x+4)+2log(x-7)-5log(x-2)+log(x^2)

OpenStudy (anonymous):

@triciaal @sexygirl7400

OpenStudy (anonymous):

i guess D

OpenStudy (anonymous):

\[\log (x^2(x^5-2x^4-71x^3-20x^2+1456x+3136))/ (x-2)^5\]

OpenStudy (anonymous):

@sammixboo @TheSmartOne

OpenStudy (anonymous):

Im stumped on this one... maybe @Michele_Laino can help

OpenStudy (zale101):

3log(x+4)+2log(x-7)-5log(x-2)+log(x^2) A number in front of logs means that they are exponents. That means, place all the numbers in front of the logs as exponents log(x+4)^(3)+log(x-7)^(2)-log(x-2)^(5)+log(x^2)

TheSmartOne (thesmartone):

if there is addition between two logs you need to make it multiplication

TheSmartOne (thesmartone):

and if there is subtraction then you need to divide it.

OpenStudy (michele_laino):

@HugoHuynh please apply this property, first: \[m \log x=\log (x ^{m})\]

TheSmartOne (thesmartone):